Question 2. (18 points) Three mutually exclusive alternative publ respective costs and benefits are included in the following table. Each of the se PW for vow years, and MARR is 15% per year, which, if any, of these projects shou lic-works projects are currently under useful life oft v0 d be selected? Use PW for YOUr calculations. You can use either B-C method. $10,500,000 $13,000,000$9,500,000 650,000 Annual Operating and Maintenance Costs Market Value 750,000 625,000 2,000,000 | 1,750,000 | 1,250,000 2,500,000 2.265,000 2,150,000 Annual Benefit
Answer #1

Project X:

Net Annual benefit = 2,500,000 - 750,000 = 1,750,000

PW = -10,500,000 + 1,750,000(P/A, 15%, 30) + 2,000,000(P/F, 15%, 30)

= -10,500,000 + 1,750,000(6.566) + 2,000,000(0.0151)

= -10,500,000 + 9,740,500 + 30,200

= -$729,300

Project Y:

Net annual benefit = 2,265,000 - 625,000 = 1,640,000

PW = -13,000,000 + 1,640,000(6.566) + 1,750,000(0.0151)

= -13,000,000 + 9,128,240 + 26,425

= -$3,845,335

Project Z:

Net annual benefit = 2,150,000 - 650,000 = 1,500,000

PW = -9,500,000 + 1,500,000(5.566) + 1,250,000(0.0151)

= -9,500,000 + 8,349,000 + 18,875

= -$1,132,125

Since PW of all the projects are negative, therefore, no projects should be selected.